OpenStudy (anonymous):

The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 38°, a = 19, c = 10

OpenStudy (anonymous):

@mathmale Can you help me? :D

OpenStudy (mathmale):

Hello! I got two help requests at the same moment, yours and someone else's, so feel obliged to divide my time between the 2 of you. Have you tried sketching this triangle? Note that we don't know yet whether its an "acute triangle" or an "obtuse triangle."

OpenStudy (mathmale):

Try to draw it "to scale."

OpenStudy (anonymous):

Okay, like this?

OpenStudy (mathmale):

Wow, oh wow! One of the nicest diagrams I've yet seen here!

OpenStudy (anonymous):

b would be like 16.16. Right?

OpenStudy (anonymous):

Thanks! I am a little confused, and I think this one might not be able to be done, because when I tried to do the SSA thing in my calculator, it said Error...

OpenStudy (mathmale):

We need to work with the data given; neither b nor B is known.

OpenStudy (anonymous):

We shouldn't try to figure those out?

OpenStudy (mathmale):

C=38 deg, c=10, a=19, A=?? Ignore B and b for now. If the triangle can be solved, then we could find B and b later. Write the Law of Sines, addressing only A, C, a and c

OpenStudy (anonymous):

When I try to do A, my calculator says error. :(

OpenStudy (mathmale):

\[\frac{ a }{ \sin A }=\frac{ c }{\sin C }\]

OpenStudy (mathmale):

Which of these variables do we know? a? c? A? C? Substitute what we know into the formula, above, which represents the "Law of Sines."

OpenStudy (anonymous):

Yup, and you get sinA=(19sin38)/10, or A=arcsin(19sin(38))/10. RIght?

OpenStudy (mathmale):

Try it. We need to know whether or not the result makes sense.

OpenStudy (anonymous):

That's what I have been trying.. I get an error thingy.

OpenStudy (mathmale):

You got an error thingy? :) ;) :)

OpenStudy (anonymous):

Oh, hush. ;p So does that mean it cannot be solved using the Law of Sines? Because that's one of my choices.. :P

OpenStudy (mathmale):

I didn't get an error thingy right off, but I did learn that sin A is supposedly 1.170. Remember, the value of the sine function MUST be between -1 and 1. So, what can you conclude?

OpenStudy (mathmale):

Please buy an error thingy and send it to me via Priority Mail.

OpenStudy (anonymous):

So it's not a sine function?

OpenStudy (anonymous):

I'll be sure to do that! :P

OpenStudy (anonymous):

A) No triangle is formed. B) A = 58.6°, B = 83.4°, b ≈ 15.6 C) A = 83.4°, B = 58.6°, b ≈ 15.6 D) The triangle cannot be solved with the Law of Sines.

OpenStudy (mathmale):

Oh, thanks!!! I wouldn't conclude that "it's not a sine function." It's just that the result you and I both got (an error thingy) is not permitted, because the value of the sine MUST be between -1 and 1. So your choice D is absolutely correct.

OpenStudy (anonymous):

Yay! Thanks! I have a few more. :(

OpenStudy (mathmale):

Go ahead and post them, one at a time. But remember, another student wanted my help at exactly the same time as u, so I have to go see whether he/she still needs it.

OpenStudy (anonymous):

Haha. You're fine! Oh, and *you. ;P